OpenStudy (anonymous):

The toddler section in a park is in the shape of a trapezoid. The parallel sides of the section measure 10 m and 14 m. The distance between the parallel sides is 8 m, as shown below. The section was remodeled to have an area that was 96 square units more than the original area. What change in the dimensions of the trapezoid was made to create the remodeled section?

OpenStudy (anonymous):

The height was multiplied by four. The length of the parallel sides and the height were doubled. The height was doubled. The length of the parallel sides and the height were multiplied by four.

OpenStudy (anonymous):

c)

OpenStudy (anonymous):

the height was doubled? how so

OpenStudy (anonymous):

okay the area of a trapezium = 1/2 (a+B) * H just plug in the values of (10=a, 14=b, 8 = H) = apparently u will get 96 Now u need to check, what could be done so that the area of the model is 96 more than its now= which means the area = "192"= 96+96 now just try those answers, the first one wouldn work, coz if u multiply the height by 4 as it is now..u will get 12*4*8= it would be sumwhere around= 384 the second one wouldnt coz then the answer would be more than 192 the third one is the correct as if u doubl the heigh= 8 *2=16, plug it n the equation and u will get 192 and thats the answer

OpenStudy (anonymous):

okay thank yoou

OpenStudy (anonymous):

No probs..just reply to my message

Directrix (directrix):

The height was doubled. ================ Original Area A = (1/2)*8* (10 + 14) A = 96. The remodeled area is 96 + 96 = 192. If the original height is doubled, the dimensions of the remodeled area are the following: h = 16, bases = 10 and 14. Remodeled Area = (1/2)* 16 * (10 + 14) Remodeled Area = 8 * 24 = 192 which is the original area plus the added area.